OpenStudy (anonymous):

PLEASE HELP ME WITH THIS ALGEBRA PROBLEM?! UGH Sarah has to figure out the area of her square garage. She knows that one side of the garage is equal to the length of her rabbit pen. The dimensions of the rectangular rabbit pen are 13 by 10. WHAT IS THE AREA OF THE GARAGE??????????

hartnn (hartnn):

what is the length of rabbit pen?

hartnn (hartnn):

hint:length is larger than width

hartnn (hartnn):

?? which is larger 13 or 10?

OpenStudy (anonymous):

13 obviously ;]

OpenStudy (anonymous):

13^2

hartnn (hartnn):

so side of garage =13 Area = side^2 = 13^2 = ?

OpenStudy (anonymous):

169

hartnn (hartnn):

yup

OpenStudy (anonymous):

so what does that mean lol is that the area

hartnn (hartnn):

AREA is 169

OpenStudy (anonymous):

AWESOME

OpenStudy (anonymous):

Two storage sheds are to have the same area. One is square and one is rectangular. The rectangular shed is 4 meters wide and 9 meters long. How large is one side of the square shed in meters????

hartnn (hartnn):

just equate the areas AREA of rect= AREA of square 4*9 = s^2 s=?

OpenStudy (anonymous):

how lol

hartnn (hartnn):

4*9=?

OpenStudy (anonymous):

area should be 169

hartnn (hartnn):

4*9=?

OpenStudy (anonymous):

36

hartnn (hartnn):

\(s^2 = 36 \\ s=\sqrt{36}=?\)

OpenStudy (anonymous):

6?

hartnn (hartnn):

thats your side!

hartnn (hartnn):

6 meters to be precise.....

OpenStudy (anonymous):

Larry wants to buy some carpeting for his living room. The length of the room is 4 times the width and the total area of the room is 16 square meters. What is the length of the living room?????

hartnn (hartnn):

lol, ask new question in new post....this is the last 1 i'll solve here l=4w AREA= l*w=16 4w*w=16 w*w=4 w=2 l=4w=8 length = 8 meters

OpenStudy (anonymous):

lol ok ;]